theyadoreshayyy:

Ellis is painting wooden fenceposts before putting them in his yard. They are each 7 feet tall and have a diameter of 1 foot. There are 11 fenceposts in all. How much paint will Ellis need to paint all the surfaces of the 11 fenceposts? Use 3.14 for π, and round your answer to the nearest hundredth. Provide an explanation and proof for your answer to receive full credit.

AZ:

Surface area of cylinder = \( 2\pi rh+2\pi r^2 \) So plug in r = 0.5 diameter is 1 so the radius is half of that and height is 7 so what is the surface area of 1 wooden fencepost?

AZ:

And then you multiply that number by 11 so you can find how much surface area you're painting for 11 fenceposts